OpenStudy (anonymous):

Complex number question: Find all the answers to (1 + sqrt(3)*i)^(1/3)

OpenStudy (anonymous):

\[\large (1 + \sqrt{3}i)^{\frac {1}{3}}\]

OpenStudy (anonymous):

\[\large (2(\cos(\frac{\pi}{3}+2\pi n)+isin(\frac{\pi}{3} + 2\pi n)))^{\frac{1}{3}}\]

OpenStudy (anonymous):